OpenStudy (anonymous):

Select the correct graph of -4x + 3y = 6.

OpenStudy (anonymous):

How do I solve this?

OpenStudy (anonymous):

can you put a picture of the graphs up, if so i can help you

OpenStudy (anonymous):

to find the x intercept,sub in 0 for y to find the y intercept,sub in 0 for x

OpenStudy (anonymous):

Alright I'll give it a try :)

OpenStudy (anonymous):

-4x + 3y = 6. so sub 0 for y... 4x + 3(0) = 6 4x = 6? /4 /4 x = 1.5?

OpenStudy (anonymous):

and sub 0 for x... -4(0) + 3y = 6 3y = 6 /3 /3 y = 2?

OpenStudy (anonymous):

or put it in y = mx + b -4x + 3y = 6 3y = 4x + 6 y = 4/3x + 2 slope(m) is 4/3. y intercept(b) = 2 plot your 2 on the y line. go up 4 and to the right 3 and up 4 and to the right 3. and you x intercept is : -4x + 3y = 6 -4x + 3(0) = 6 -4x = 6 x = -6/4 x = - 1.5

OpenStudy (anonymous):

plot your point on the x line at -1.5 and on the y line at 2.

OpenStudy (anonymous):

and from the -1.5, go up 4 and to the right 3 (rise/run) and it should cross the y line at 2

OpenStudy (anonymous):

so the location on the chart would be -1.4, 2 and the slope 4/3?

OpenStudy (anonymous):

-1.5 is x intercept 2 is y intercept slope is 4/3

OpenStudy (anonymous):

Could you explain to me how you found the slope a bit more?

OpenStudy (anonymous):

would that be the line?

OpenStudy (anonymous):

I think it is because the line intercepts x at 1.5 and 2 at the y line. and the slope is also up 4 over 3

OpenStudy (anonymous):

of course....in the formula y = mx + b, the slope is in the m spot. so put your equation in the y = mx + b form. -4x + 3y = 6 (add 4 to both sides) 3y = 4x + 6 (now divide by 3) y = 4/3x + 2 you can see the m spot has 4/3, so that is your slope. you are correct.....that is the line

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

I think I'm good, thanks for your help :)

OpenStudy (anonymous):

I'm probably going to be posting some more questions like this aswell :)